# A SAM-guided Two-stream Lightweight Model for Anomaly Detection

### v1

@@ 3.2.2. The Decoder
Inspired by (Zhang et al., 2023a), the image embeddings generated by the student encoder can closely approximate that of the original teacher encoder. This observation leads to the conclusion that the utilization of distinct decoders for the TLM may not be imperative. Instead, a shared two-layer decoder as illustrated in Figure 4, is an efficient alternative. We observe that the shared decoder, while economical in terms of parameters and computation time, achieves a satisfactory result with a slight decrease compared to employing two separated decoders, shown in Figure 5.
@@ 3.2.3. Remark
The difference between our STLM and DeSTSeg (Zhang et al., 2023c) lies in 1) our method learns not only a denoising student but also a plain student in our TLM. The plain student stream can learn generalized knowledge related to the anomaly detection task, which enables it to effectively represent normal features and even previously unseen patterns in the training set, ensuring differences between features of two streams are well-captured. 2) The high-generalization-capable SAM, as the teacher network of our TLM, also helps the denoising student to generate high-quality reconstructed results for genuine anomalies.
Figure 6. Comparison between lightweight feature aggregation module and original feature aggregation module.
Besides, it is worth noting that our method exhibits superior performance and a more compact model compared to DeSTSeg.
@@ 3.3. Feature Aggregation Module
Previous studies (Salehi et al., 2021; Wang et al., 2021) have noted that sub-optimal results can occur when distinctions among features at different levels are not uniformly precise. Following an extensive series of experiments, it has become evident that the inclusion of a segmentation network, guiding feature fusion through additional supervision signals, improves performances. The feature aggregation (FA) module is composed of two residual blocks and an atrous spatial pyramid pooling (ASPP) module (Chen et al., 2017). As a primary consideration, we have refrained from constraining the weights of the TLM when training the segmentation network.
Notably, our findings indicate that simultaneous training of these components yields enhanced results. Furthermore, it has been observed that reducing the depth of the FA module does not significantly impact the performance as indicated in Figure 6. As a result, an effort has been made to make this module more lightweight by reducing the channels from 256 to 128 and adjusting the dilation rate to [1, 1, 3], following a similar method in (Zhang et al., 2023e). It mitigates memory costs associated with both training and inference, which is a factor of paramount importance in practical implementations.
When training, synthetic anomalous images are employed as inputs for the TLM, with the corresponding binary anomaly mask serving as the ground truth. The decoder generates features that have the same shape as the ground truth mask $M$. The similarities are calculated using Equation (2) and then concatenated as $\hat{X}$, which is subsequently fed into the FA module. Inspired by (Zavrtanik et al., 2021a; Yang et al., 2023), a focal loss (Lin et al., 2017) and an L1 loss (Girshick, 2015) are applied to increase the robustness toward accurate segmentation of challenging examples and reduce over-sensitivity to outliers, respectively. The FA module outputs an anomaly score map $M_{ij}^{o}$, which is of the same shape as the ground truth mask $M$.
@@
where the focus ($\gamma$) is set as 4, following the set in (Zhang et al., 2023d, c).
@@ 3.4. Training and Inference
For inference, the fixed SAM is discarded and the procedure is presented in Figure 2 (b). For pixel $ij$, the pixel-level anomaly segmentation map $M_{ij}$ is provided by the end of the network. It is anticipated that the output will have higher values for pixels that are anomalous pixels. For the computation of the image-level anomaly score, we take the average of the top-$K$ anomalous pixel values from the anomaly score map, following (Zhang et al., 2023d, c).
@@ 4.1.1. Datasets
Additionally, we extend our experimentation on the VisA (Zou et al., 2022) and DAGM (Wieler and Hahn, 2007) dataset to showcase the generalization capabilities of STLM facing more complex datasets with. The VisA Dataset comprises 10,821 high-resolution color images (9,621 normal and 1,200 anomalous samples) encompassing 12 objects across 3 domains, establishing it as the most extensive anomaly detection dataset in the industrial domain to date. DAGM contains 10 textured objects with small abnormal regions that bear a strong visual resemblance to the background. For each class, we first move all anomalous samples from the original training set to the original test set and then move all normal samples from the test set to the training set. Furthermore, we randomly select 30 normal samples from the training set, designating them as ”good” samples in the test set, which are subsequently removed from the training set.
Additionally, we use the Describable Textures Dataset (DTD) (Cimpoi et al., 2014) as the source of anomaly images (denoted as A in Equation (1)).
@@ 4.1.2. Evaluation Metrics
In line with previous research, we employ AUC (Area Under the ROC Curve) to evaluate image-level and pixel-level anomaly detection. However, anomalous regions typically only occupy a tiny fraction of the entire image. Consequently, Pixel-AUROC may not accurately reflect the localization accuracy, as the false positive rate is primarily influenced by the vast number of anomalous-free pixels (Tao et al., 2022). To offer a more comprehensive measure of localization performance, we introduce the Per Region Overlap (PRO) (Bergmann et al., 2020), which assigns equal weight to anomaly regions of varying sizes. Pixel-level Average Precision (AP) (Saito and Rehmsmeier, 2015) is also introduced for comprehensive measuring. The PRO score treats anomaly regions of varying sizes equally, whereas AP is more suitable for highly imbalanced classes, particularly in the context of industrial anomaly localization where accuracy plays a critical role.
@@
Table 2. Anomaly Detection and Localization on MVTec (Bergmann et al., 2019a), VisA (Zou et al., 2022) and DAGM (Zavrtanik et al., 2021a)
@@
Method DRAEM (Zavrtanik et al., 2021a) CFLOW (Gudovskiy et al., 2022) PatchCore (Roth et al., 2022) RD4AD (Deng and Li, 2022) SimpleNet (Liu et al., 2023) DeSTSeg (Zhang et al., 2023c) RD++ (Tien et al., 2023) FOD (Yao et al., 2023a) Ours
Carpet 96.90/97.50 97.60/99.20 99.10/99.00 98.70/98.90 99.70/98.50 -/98.30 100/99.20 100/- 99.48/99.91
Grid 99.90/99.70 98.10/98.90 97.30/98.70 100/98.30 99.70/98.80 -/99.20 100/99.30 100/- 95.57/95.40
Leather 100/99.00 99.90/99.70 100/99.30 100/99.40 100/99.20 -/99.70 100/99.40 100/- 100/99.11
Tile 100/99.20 97.10/96.20 99.30/95.80 99.70/95.70 99.80/97.00 -/99.10 99.70/96.60 100/- 100/99.58
Wood 99.50/95.50 98.70/86.00 99.60/95.10 99.50/95.80 100/94.50 -/98.00 99.30/95.80 99.1/- 100/97.99
Bottle 98.00/99.10 99.90/97.20 100/98.60 100/98.80 100/98.00 -/99.40 100/98.80 100/- 100/97.77
Cable 90.90/95.20 97.60/97.80 99.90/98.50 96.10/97.20 99.90/97.60 -/97.70 99.20/98.40 99.50/- 98.98/96.95
Capsule 91.30/88.10 97.00/99.10 98.00/99.00 96.10/98.70 97.70/98.90 -/99.10 99.00/98.80 100.00/- 98.69/98.48
Hazelnut 100/99.70 100/98.80 100/98.70 100/99.00 100/97.90 -/99.80 100/99.20 100/- 100/ 99.69
Metal_nut 100/99.60 98.50/98.60 99.90/98.30 100/97.30 100/98.80 -/99.00 100/98.10 100/- 100/99.45
Pill 97.10/97.30 96.20/98.90 97.50/97.60 98.70/98.10 99.00/98.60 -/99.10 98.40/98.30 98.40/- 98.20/98.61
Screw 98.70/99.30 93.10/98.90 98.20/99.50 97.80/99.70 98.20/99.30 -/98.80 98.90/99.70 96.7/- 99.47/97.56
Toothbrush 100/97.30 98.80/99.00 100/98.60 100/99.10 99.70/98.50 -/99.40 100/99.10 94.4/- 99.06/99.61
Transistor 91.70/85.20 92.90/98.20 99.90/96.50 95.50/92.30 100/97.60 -/92.50 98.50/94.30 100.0/- 97.58/94.81
Zipper 100/99.10 97.10/99.10 99.50/98.90 97.90/98.30 99.90/98.90 -/99.60 98.60/98.80 99.70/- 99.74/98.93
Average 97.60/96.70 97.50/97.70 99.20/98.10 98.70/97.80 99.57/98.14 99.00/98.20 99.44/98.25 99.20/98.30 99.05/98.26
Inference time 159 37 180 28 39 18 30 103 20
Parameters 97 94.7 186.55 150.64 52.88 35.16 154.87 28.83 16.56
Method VisA MVTec DAGM Average
I ↑ P ↑ O ↑ A ↑ I ↑ P ↑ O ↑ A ↑ I ↑ P ↑ O ↑ A ↑ I ↑ P ↑ O ↑ A ↑
PatchCore (Li et al., 2021b) (186.6M) 95.10 98.80 91.20 40.10 99.20 98.10 93.40 56.10 93.60 96.70 89.30 51.70 95.97 97.87 91.30 49.30
RD4AD (Roth et al., 2022) (150.6M) 96.00 90.10 70.90 27.70 98.70 97.80 93.93 58.00 95.80 97.50 93.00 53.40 97.95 98.12 94.06 55.30
RD++ (Tien et al., 2023) (154.9M) 95.90 98.70 93.40 40.80 99.44 98.25 94.99 60.80 98.50 97.40 93.80 64.30 96.70 98.20 94.30 52.40
DeSTSeg (Zhang et al., 2023c) (35.2M) 91.95 97.73 90.02 40.76 99.00 98.20 95.11 75.80 97.44 93.55 87.88 56.99 96.13 96.49 91.00 57.85
SimpleNet (Liu et al., 2023) (52.9M) 96.80 97.80 88.70 36.30 99.57 98.14 90.00 54.80 95.30 97.10 91.30 48.10 97.22 97.68 90.00 46.40
FastFlow (92M) 82.20 88.20 59.80 15.60 90.50 95.50 85.60 39.80 87.40 91.10 79.90 34.20 86.70 91.60 75.10 29.87
DRAEM (Zavrtanik et al., 2021a) (97M) 88.70 94.40 73.70 30.50 97.60 96.70 92.10 68.40 90.80 86.80 71.00 30.60 92.37 92.63 78.93 43.17
Ours (16.6M) 96.73 98.36 93.83 47.63 99.05 98.26 94.92 76.32 98.30 96.33 91.14 64.91 98.03 97.65 93.30 62.95
“I”, “P”, “O” and “A” respectively refer to the five metrics of image-level AUROC, pixel-level AUROC, PRO and AP. The best results on PRO are highlighted in bold.
Table 2. Anomaly Detection and Localization on MVTec (Bergmann et al., 2019a), VisA (Zou et al., 2022) and DAGM (Zavrtanik et al., 2021a)
Our comparison SOTA methods include DRAEM (Zavrtanik et al., 2021a), CFLOW (Gudovskiy et al., 2022), CFA (Lee et al., 2022), PatchCore (Roth et al., 2022), RD4AD (Deng and Li, 2022), SimpleNet (Liu et al., 2023), DeSTSeg (Zhang et al., 2023c), RD++ (Tien et al., 2023), FastFlow (Yu et al., 2021), and FOD (Yao et al., 2023a).
Figure 8. Visualization examples of our method on VisA.
4.1.3. Implementation Details
Our initialized two-stream encoders of TLM are the ViT-Tiny (Wu et al., 2022) with MoblieSAM (Zhang et al., 2023a) weight, while SAM (Kirillov et al., 2023) serves as the teacher network. All images in the two datasets are resized to 1024 × 1024. We employ the Adam Optimizer (Kingma and Ba, 2014), setting the learning rate at 0.0005 for the TLM. For the FA module, we choose the Stochastic Gradient Descent Optimizer, using the same configuration as presented in (Zhang et al., 2023c). Each loss function in our paper carries equal weight. We conduct training for 200 epochs with a batch size of 2 and compute the average of the top 100 anomalous pixels as the image-level anomaly score. Notably, we use the augmentation method, i.e., pseudo anomaly generation strategy, that most state-of-the-art methods use for fair competition. Prototypical (Zhang et al., 2023d) presents more complex anomaly generation strategies to get better performance.
[…]
Class CFA (Lee et al., 2022) PatchCore (Roth et al., 2022) RD4AD (Deng and Li, 2022) DeSTSeg (Zhang et al., 2023c) RD++ (Tien et al., 2023) Ours
Carpet 96.54 96.60 97.00 97.12 97.70 96.97
Grid 94.04 96.00 97.60 96.40 97.70 91.56
Leather 97.43 98.90 99.10 98.88 99.20 98.05
Tile 89.26 87.30 90.60 97.63 92.40 99.23
Wood 90.54 89.40 90.90 95.26 93.30 98.41
Avg. Text. 93.56 93.64 95.04 97.06 96.06 96.84
Bottle 95.76 96.20 96.60 97.09 97.00 90.92
Cable 94.17 92.50 91.00 83.19 93.90 92.11
Capsule 93.66 95.50 95.80 95.12 96.40 97.28
Hazelnut 95.75 93.80 95.50 98.09 96.30 98.29
Metal_nut 94.54 91.40 92.30 94.54 93.00 92.95
Pill 97.19 93.20 96.40 95.12 97.00 94.69
Screw 95.23 97.90 98.20 93.66 98.60 98.60
Toothbrush 91.14 91.50 94.50 96.22 94.20 96.14
Transistor 95.35 83.70 78.00 90.94 81.80 82.50
Zipper 95.95 97.10 95.40 97.35 96.30 96.15
Avg. Obj. 94.87 93.28 93.37 94.13 94.45 93.96
Average 94.44 93.4 93.93 95.11 94.99 94.92
Parameters 25.6M 186.6M 150.6M 35.1M 154.9M 16.6M
@@ 4.2.1. Anomaly Detection and Localization on MVTec
For the purpose of ensuring fair comparisons with other studies, we use the results reported in the original papers or the results re-evaluated by (Zhang et al., 2023c, d). In cases where results are not available, we indicate with a hyphen (‘-’).
Table 2 presents the results of anomaly detection and localization on the MVTec dataset. On average, our method achieves a competitive performance and secures the highest score for 9 out of 15 classes. To provide a comprehensive overview of anomaly localization capabilities, we include additional PRO metric results in Table 3. In 5 out of 15 classes, our method achieves a superior PRO score and it remains comparable to state-of-the-art in the remaining classes. The specific results in Figure 1 are shown in Table 2, our method achieves the most efficient method and the second-fastest inference speed, demonstrating the model efficiency and mobile-friendliness of our STLM.
All experiments are consistently conducted on an NVIDIA GeForce RTX 3090. Notably, we attempt to analyze the reason why our STLM has fewer parameters than DeSTSeg but longer inference times is the software and hardware optimization works for ResNet-like frameworks are more mature compared to Transformer (Li et al., 2023). The optimal inference speed of the image encoder requires further study, which we plan to optimize in future work.

### v2

@@ 3.2.2. The Decoder
Inspired by (Zhang et al., 2023a), the image embeddings generated by the student encoder can closely approximate that of the original teacher encoder. This observation leads to the conclusion that the utilization of distinct decoders for the TLM may not be imperative. Instead, a shared two-layer decoder, is an efficient alternative. We observe that the shared decoder, while economical in terms of parameters and computation time, achieves a satisfactory result with a slight decrease compared to employing two separated decoders, shown in Figure 7(a).
@@ 3.2.3. Remark
The difference between our STLM and DeSTSeg (Zhang et al., 2023b) lies in 1) our method learns not only a denoising student but also a plain student in our TLM. The plain student stream can learn generalized knowledge related to the anomaly detection task, which enables it to effectively represent normal features and even previously unseen patterns in the training set, ensuring differences between features of two streams are well-captured. 2) The high-generalization-capable SAM, as the teacher network of our TLM, also helps the denoising student to generate high-quality reconstructed results for genuine anomalies. Besides, it is worth noting that our method exhibits superior performance and a more compact model compared to DeSTSeg.
@@ 3.3. Feature Aggregation Module
Previous studies (Salehi et al., 2021; Wang et al., 2021) have noted that sub-optimal results can occur when distinctions among features at different levels are not uniformly precise. Following an extensive series of experiments, it has become evident that the inclusion of a segmentation network, guiding feature fusion through additional supervision signals, improves performances. The feature aggregation (FA) module is composed of two residual blocks and an atrous spatial pyramid pooling (ASPP) module. As a primary consideration, we have refrained from constraining the weights of the TLM when training the segmentation network.
Notably, our findings indicate that simultaneous training of these components yields enhanced results. Furthermore, it has been observed that reducing the depth of the FA module does not significantly impact the performance as indicated in Figure 7(b). As a result, an effort has been made to make this module more lightweight by reducing the channels from 256 to 128 and adjusting the dilation rate to [1, 1, 3], following a similar method in (Zhang et al., 2023d). It mitigates memory costs associated with both training and inference, which is a factor of paramount importance in practical implementations.
When training, synthetic anomalous images are employed as inputs for the TLM, with the corresponding binary anomaly mask serving as the ground truth. The decoder generates features that have the same shape as the ground truth mask $M$. The similarities are calculated using Equation (2) and then concatenated as $\hat{X}$, which is subsequently fed into the FA module, shown in Figure 2 (d). Inspired by (Zavrtanik et al., 2021a), a focal loss and an L1 loss are applied to increase the robustness toward accurate segmentation of challenging examples and reduce over-sensitivity to outliers, respectively. The FA module outputs an anomaly score map $M_{ij}^{o}$, which is of the same shape as the ground truth mask $M$.
@@
where the focus ($\gamma$) is set as 4, following the set in (Zhang et al., 2023c, b).
@@ 3.4. Training and Inference
For inference, the fixed SAM is discarded and the procedure is presented in Figure 2 (b). For pixel $ij$, the pixel-level anomaly segmentation map $M_{ij}$ is provided by the end of the network. It is anticipated that the output will have higher values for pixels that are anomalous pixels. For the computation of the image-level anomaly score, we take the average of the top-$K$ anomalous pixel values from the anomaly score map, following (Zhang et al., 2023c, b).
Figure 3. Visualization examples of VisA (Zou et al., 2022), MVTec LOCO (Bergmann et al., 2022), DAGM (Zavrtanik et al., 2021a) and MVTec (Bergmann et al., 2019a).
@@ 4.1.1. Datasets
Additionally, we extend our experimentation on the VisA (Zou et al., 2022), MVTec LOCO (Bergmann et al., 2022) and DAGM (Wieler and Hahn, 2007) dataset to showcase the generalization capabilities of STLM facing more complex datasets with, examples in Figure 3. The VisA Dataset comprises 10,821 high-resolution color images (9,621 normal and 1,200 anomalous samples) encompassing 12 objects across 3 domains, establishing it as the most extensive anomaly detection dataset in the industrial domain to date. The MVTec LOCO dataset includes both structural and logical anomalies. It contains 3644 images from five different categories inspired by real-world industrial inspection scenarios. Structural anomalies appear as scratches, dents, or contaminations in the manufactured products. Logical anomalies violate underlying constraints, e.g., a permissible object being present in an invalid location or a required object not being present at all. Notably, for the convenience of comparison, we only select structural anomalies. DAGM contains 10 textured objects with small abnormal regions that bear a strong visual resemblance to the background. For each class, we first move all anomalous samples from the original training set to the original test set and then move all normal samples from the test set to the training set. Furthermore, we randomly select 30 normal samples from the training set, designating them as ”good” samples in the test set.
@@ 4.1.2. Evaluation Metrics
In line with previous research, we employ AUC (Area Under the ROC Curve) to evaluate image-level and pixel-level anomaly detection. However, anomalous regions typically only occupy a tiny fraction of the entire image. Consequently, Pixel-AUROC may not accurately reflect the localization accuracy, as the false positive rate is primarily influenced by the vast number of anomalous-free pixels (Tao et al., 2022). To offer a more comprehensive measure of localization performance, we introduce the Per Region Overlap (PRO) (Bergmann et al., 2020), which assigns equal weight to anomaly regions of varying sizes. Pixel-level Average Precision (AP) is also introduced for comprehensive measuring. The PRO score treats anomaly regions of varying sizes equally, whereas AP is more suitable for highly imbalanced classes, particularly in the context of industrial anomaly localization where accuracy plays a critical role. Our comparison SOTA methods include DRAEM (Zavrtanik et al., 2021a), CFLOW (Gudovskiy et al., 2022), CFA (Lee et al., 2022), PatchCore (Roth et al., 2022), RD4AD (Deng and Li, 2022), SimpleNet (Liu et al., 2023), DeSTSeg (Zhang et al., 2023b), RD++ (Tien et al., 2023), FastFlow (Yu et al., 2021), and FOD (Yao et al., 2023a).
4.1.3. Implementation Details
Our initialized two-stream encoders of TLM are the ViT-Tiny (Wu et al., 2022) with MoblieSAM weight, while SAM serves as the teacher network. All images in the two datasets are resized to 1024 × 1024. We employ the Adam Optimizer, setting the learning rate at 0.0005 for the TLM. For the FA module, we choose the Stochastic Gradient Descent Optimizer, using the same configuration as presented in (Zhang et al., 2023b). Each loss function in our paper carries equal weight. We conduct training for 200 epochs with a batch size of 2 and compute the average of the top 100 anomalous pixels as the image-level anomaly score. Notably, we use the augmentation method, i.e., pseudo anomaly generation strategy, that most state-of-the-art methods use for fair competition. Prototypical (Zhang et al., 2023c) presents more complex anomaly generation strategies to performe better.
@@ 4.1.3. Implementation Details
      Method DRAEM (Zavrtanik et al., 2021a) CFLOW (Gudovskiy et al., 2022) PatchCore (Roth et al., 2022) RD4AD (Deng and Li, 2022) Ours
Carpet 96.90/97.50 97.60/99.20 99.10/99.00 98.70/98.90 99.48/99.91
Grid 99.90/99.70 98.10/98.90 97.30/98.70 100/98.30 95.57/95.40
Leather 100/99.00 99.90/99.70 100/99.30 100/99.40 100/99.11
Tile 100/99.20 97.10/96.20 99.30/95.80 99.70/95.70 100/99.58
Wood 99.50/95.50 98.70/86.00 99.60/95.10 99.50/95.80 100/97.99
Bottle 98.00/99.10 99.90/97.20 100/98.60 100/98.80 100/97.77
Cable 90.90/95.20 97.60/97.80 99.90/98.50 96.10/97.20 98.98/96.95
Capsule 91.30/88.10 97.00/99.10 98.00/99.00 96.10/98.70 98.69/98.48
Hazelnut 100/99.70 100/98.80 100/98.70 100/99.00 100/ 99.69
Metal_nut 100/99.60 98.50/98.60 99.90/98.30 100/97.30 100/99.45
Pill 97.10/97.30 96.20/98.90 97.50/97.60 98.70/98.10 98.20/98.61
Screw 98.70/99.30 93.10/98.90 98.20/99.50 97.80/99.70 99.47/97.56
Toothbrush 100/97.30 98.80/99.00 100/98.60 100/99.10 99.06/99.61
Transistor 91.70/85.20 92.90/98.20 99.90/96.50 95.50/92.30 97.58/94.81
Zipper 100/99.10 97.10/99.10 99.50/98.90 97.90/98.30 99.74/98.93
Average 97.60/96.70 97.50/97.70 99.20/98.10 98.70/97.80 99.05/98.26
Inference time 159 37 180 28 20
Parameters 97 94.7 186.55 150.64 16.56
Table 2. Table 1 (continued)
      Method SimpleNet (Liu et al., 2023) DeSTSeg (Zhang et al., 2023b) RD++ (Tien et al., 2023) FOD (Yao et al., 2023a) Ours
Carpet 99.70/98.50 -/98.30 100/99.20 100/- 99.48/99.91
Grid 99.70/98.80 -/99.20 100/99.30 100/- 95.57/95.40
Leather 100/99.20 -/99.70 100/99.40 100/- 100/99.11
Tile 99.80/97.00 -/99.10 99.70/96.60 100/- 100/99.58
Wood 100/94.50 -/98.00 99.30/95.80 99.1/- 100/97.99
Bottle 100/98.00 -/99.40 100/98.80 100/- 100/97.77
Cable 99.90/97.60 -/97.70 99.20/98.40 99.50/- 98.98/96.95
Capsule 97.70/98.90 -/99.10 99.00/98.80 100.00/- 98.69/98.48
Hazelnut 100/97.90 -/99.80 100/99.20 100/- 100/ 99.69
Metal_nut 100/98.80 -/99.00 100/98.10 100/- 100/99.45
Pill 99.00/98.60 -/99.10 98.40/98.30 98.40/- 98.20/98.61
Screw 98.20/99.30 -/98.80 98.90/99.70 96.7/- 99.47/97.56
Toothbrush 99.70/98.50 -/99.40 100/99.10 94.4/- 99.06/99.61
Transistor 100/97.60 -/92.50 98.50/94.30 100.0/- 97.58/94.81
Zipper 99.90/98.90 -/99.60 98.60/98.80 99.70/- 99.74/98.93
Average 99.57/98.14 99.00/98.20 99.44/98.25 99.20/98.30 99.05/98.26
Inference time 39 18 30 103 20
Parameters 52.88 35.16 154.87 28.83 16.56
@@ 4.2.1. Anomaly Detection and Localization on MVTec
For the purpose of ensuring fair comparisons with other studies, we use the results reported in the original papers or the results re-evaluated by (Zhang et al., 2023b, c). In cases where results are not available, we indicate with a hyphen (‘-’).
Table 1 presents the results of anomaly detection and localization on the MVTec dataset. On average, our method achieves a competitive performance and secures the highest score for 9 out of 15 classes. The specific results in Figure 1 are shown in Table 1, our method achieves the most efficient method and the second-fastest inference speed, demonstrating the model efficiency and mobile-friendliness of our STLM.
All experiments are consistently conducted on an NVIDIA GeForce RTX 3090. Notably, we attempt to analyze the reason why our STLM has fewer parameters than DeSTSeg but longer inference times is the software and hardware optimization works for ResNet-like frameworks are more mature compared to Transformer (Li et al., 2023a). The optimal inference speed of the image encoder requires further study, which we plan to optimize in future work.
Table 3. Anomaly Detection and Localization on MVTec (Bergmann et al., 2019a), VisA (Zou et al., 2022), MVTec LOCO (Bergmann et al., 2022) and DAGM (Zavrtanik et al., 2021a)
            Method VisA MVTec
I ↑ P ↑ O ↑ A ↑ I ↑ P ↑ O ↑ A ↑
PatchCore (Li et al., 2021b) (186.6M) 95.10 98.80 91.20 40.10 99.20 98.10 93.40 56.10
RD4AD (Roth et al., 2022) (150.6M) 96.00 90.10 70.90 27.70 98.70 97.80 93.93 58.00
RD++ (Tien et al., 2023) (154.9M) 95.90 98.70 93.40 40.80 99.44 98.25 94.99 60.80
DeSTSeg (Zhang et al., 2023b) (35.2M) 90.02 40.76 99.00 98.20 95.11 75.80 89.72 57.85
SimpleNet (Liu et al., 2023) (52.9M) 88.70 36.30 99.57 98.14 90.00 54.80 82.65 46.40
FastFlow (Yu et al., 2021) (92M) 59.80 15.60 90.50 95.50 85.60 39.80 77.38 29.87
DRAEM (Zavrtanik et al., 2021a) (97M) 73.70 30.50 97.60 96.70 92.10 68.40 74.40 43.17
Ours (16.6M) 93.83 47.63 99.05 98.26 94.92 76.32 93.18 90.51
Table 4. Table 3 (continued)
  Method MVTec LOCO DAGM Average
I ↑ O ↑ I ↑ P ↑ O ↑ A ↑ I ↑ P ↑ O ↑ A ↑
PatchCore 84.80 64.30 93.60 96.70 89.30 51.70 93.18 97.87 84.55 49.30
RD4AD 88.00 84.80 95.80 97.50 93.00 53.40 94.63 95.13 85.66 46.37
RD++ 92.45 55.30 98.50 97.40 93.80 64.30 95.41 98.12 92.45 55.30
DeSTSeg 91.95 97.73 97.44 93.55 87.88 56.99 94.71 96.49 89.72 57.85
SimpleNet 96.80 97.80 95.30 97.10 91.30 48.10 93.84 97.68 82.65 46.40
FastFlow 82.20 88.20 87.40 91.10 79.90 34.20 85.75 91.60 77.38 29.87
DRAEM 88.70 94.40 90.80 86.80 71.00 30.60 87.88 92.63 74.40 43.17
Ours 96.73 98.36 98.30 96.33 91.14 64.91 96.81 97.65 92.60 62.95
“I”, “P”, “O” and “A” respectively refer to the five metrics of image-level AUROC, pixel-level AUROC, PRO and AP. The best results are highlighted in bold.
Table 5. Additional experiments on model parameters and inference flexibility.
        Method Inference time (ms) Parameter (×$\times 10^{6}$) GFLOPS (×$\times 10^{9}$) GPU Memory (MiB)
DRAEM (Zavrtanik et al., 2021a) 159 97.4 696 -
FastFlow (Yu et al., 2021) 43 92.18 108 6348
PatchCore (Roth et al., 2022) 180 186.55 66+kNN 10010+kNN
RD4AD (Deng and Li, 2022) 28 150.64 39 3219
SimpleNet (Liu et al., 2023) 39 52.88 57 7983
DeSTSeg (Zhang et al., 2023b) 18 35.16 40 2527
RD++ (Tien et al., 2023) 30 166.10 57 4761
STLM (Ours) 20 16.56 55 2005
Figure 5. Visualization examples of our method on VisA.